40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 10 22:56:17 2011, kiwamu, Update, SUS, f2p filters installed on ETMY f2p_etmy.png
    Reply  Wed May 11 11:44:52 2011, kiwamu, Update, SUS, f2p filters installed on ITMY f2p_itmy.png
Message ID: 4688     Entry time: Wed May 11 11:44:52 2011     In reply to: 4682
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: f2p filters installed on ITMY 

New f2p filters were installed on ITMY this morning. The statistical error of the coil gain setttigs are about 0.6 % at high frequency.

NEXT : PRM, SRM, BS, ITMX and ETMX.

 

f2p_itmy.png

 Pendulum mode = 0.988 Hz, Q = 1

C1:SUS-ITMY_ULCOIL_GAIN = 1.02482

C1:SUS-ITMY_URCOIL_GAIN = -1.06831

C1:SUS-ITMY_LLCOIL_GAIN = -0.996671

C1:SUS-ITMY_LRCOIL_GAIN = 0.91079

UL: fz = 1.014824 Hz, Q = 1.027150

UR: fz = 0.975038 Hz, Q = 0.98688

LL: fz = 1.000229 Hz, Q = 1.012378

LR: fz = 0.972688 Hz, Q = 0.946116

Quote from #4682

New f2p filters were installed on ETMY. 

The coil imbalances are now about 0.8% at high frequency (i.e. above the resonant freq of the pendulum mode)

ELOG V3.1.3-